a company uses 3 machines A,B,C.propotion of products produced by the machine are .2,.48,.22 . the percent defective for each machines is 1% 2% and 3% of their individual output. a defective output is drawn from a days production.what is probability that it was not prpduced by A?.
1
Expert's answer
2019-12-03T09:48:08-0500

Suppose E is the event of getting defected output.

Then from the question, we have:

"P(A)=0.2,P(B)=0.48,P(C)=0.22"

"P(E\/A)=0.01,P(E\/B)=0.02,P(E\/C)=0.03"

Now, we need to find the probability that if the bulb is defective then it is not produced by A, that means it is produced by either B or C.

That is, we need to find "P(B\/E)+P(C\/E)"

Now,

"P(B\/E)=(P(B)*P(E\/B))\/(P(A)*P(E\/A)+P(B)*P(E\/B)+P(C)*P(E\/C))"

"P(B\/E)=(0.48*0.02)\/(0.2*0.01+0.48*0.02+0.22*0.03)"

"=0.0096\/(0.0020+0.0096+0.0066)"

"=0.0096\/0.0182"

"=0.5275"


Now, similarly:

"P(C\/E)=(P(C)*P(E\/C))\/(P(A)*P(E\/A)+P(B)*P(E\/B)+P(C)*P(E\/C))"

"P(C\/E)=(0.22*0.03)\/(0.2*0.01+0.48*0.02+0.22*0.03)"

"=0.0066\/(0.0020+0.0096+0.0066)"

"=0.0066\/0.0182"

"=0.3626"


So, the probability that the if the output is defected then it was not produced by A will be equal to ,

"P(B\/E)+P(C\/E)=0.5275+0.3626"

"=0.8901"